OpenStudy (anonymous):

will give a medal! Find the angle the normal makes with a positive x-axis 3x-2y-1=0 5x+y-12=0 4x+3y-4=0

OpenStudy (anonymous):

3x-2y-1=0 2y=3x-1 \[y=\frac{ 3 }{ 2 }x-\frac{ 1 }{ 2 }\] slope of line=3/2 slope of normal to the line=-1/m=-2/3

OpenStudy (anonymous):

so how do you work out the angle from that

OpenStudy (anonymous):

@surjithayer